### v1

# Adversarial Agents: Black-Box Evasion Attacks with Reinforcement Learning

Abstract
Reinforcement learning (RL) offers powerful techniques for solving complex sequential decision-making tasks from experience. In this paper, we demonstrate how RL can be applied to adversarial machine learning (AML) to develop a new class of attacks that learn to generate adversarial examples: inputs designed to fool machine learning models. Unlike traditional AML methods that craft adversarial examples independently, our RL-based approach retains and exploits past attack experience to improve future attacks. We formulate adversarial example generation as a Markov Decision Process and evaluate RL’s ability to (a) learn effective and efficient attack strategies and (b) compete with state-of-the-art AML. On CIFAR-10, our agent increases the success rate of adversarial examples by 19.4% and decreases the median number of victim model queries per adversarial example by 53.2% from the start to the end of training. In a head-to-head comparison with a state-of-the-art image attack, SquareAttack, our approach enables an adversary to generate adversarial examples with 13.1% more success after 5000 episodes of training. From a security perspective, this work demonstrates a powerful new attack vector that uses RL to attack ML models efficiently and at scale.
1 Introduction
The advancement of AI has led to an explosion of applications that rely on the decision-making and generation capabilities of ML models. Such models have revolutionized online assistants, commerce, content generation, cybersecurity, entertainment and gaming, to name just a few. At the same time, adversaries wishing to manipulate these applications have developed algorithms that produce inputs that are intended to fool the model, e.g., bypass a content filter. The related field–called adversarial machine learning (AML)–studies algorithms for adversarial examples targeted at victim models (Papernot et al., 2016). Extant AML uses optimization algorithms to minimize the distortion (called perturbations) applied to an input such that the resultant input with distortion is misclassified by the victim ML model (Madry et al., 2019; Carlini & Wagner, 2017). However, the technical community’s understanding of model defenses and adversarial capabilities is at best limited.
Adversarial example generation algorithms rely on gradient-based optimization that operate in isolation from any other information (Madry et al., 2019; Carlini & Wagner, 2017; Andriushchenko et al., 2020). These approaches fail to leverage information gained from previous attacks to improve subsequent attacks on other data. This presents a gap in AML that studies if adversarial examples can be learned, i.e., if the effectiveness and efficiency of the attack improves with experience. If an adversary with restricted access (called black-box access) to a victim model were able to produce adversarial examples at scale (e.g., distributed denial of service attack), this would represent a major new capability for adversaries attacking machine learning.
We posit that modeling an AML adversary as an RL agent will enable attack strategies that become increasingly more efficient and effective over time–which is a capability not currently available in existing work. In this paper, we introduce and evaluate an attack based on RL to generate adversarial examples. Modeled as an RL agent, the adversary learns which perturbations are the most effective at fooling the model. Once a policy is learned, the adversary then uses the policy to generate adversarial examples. Thus, an adversarial agent can attack a model without a costly gradient optimization.
Figure 1: Overview of the reinforcement learning attack with CIFAR-10. The adversary interacts with the victim model by querying images and receiving feedback, iteratively generating adversarial examples while adapting its strategy based on new input data.
We demonstrate that adversarial example generation can be modeled as a Markov Decision Process (MDP). The MDP formulation allows us to readily use RL in such a way that encapsulates the underlying semantics of the attack: input samples and victim model outputs as states, perturbations as actions, and differences in adversarial objectives as rewards. We introduce two RL attacks that reflect the two categories of traditional AML algorithms: RL Max Loss and RL Min Norm. The attacks use an RL algorithm to train the agent, then are used in a policy evaluation setting to craft adversarial examples. The attacks are performed on the CIFAR-10 image classification task against a ResNet50 (He et al., 2015) victim model with the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017). Under this framework, we perform a multi-step evaluation of the agent with respect to (a) learning, (b) tuning and (c) accuracy relative to known approaches.
First, we evaluate if an RL agent can learn adversarial examples. In other words, can RL improve the effectiveness and efficiency of adversarial examples over training? Indeed, both RL Max Loss and RL Min Norm attacks increase average return over training, validating RL’s ability to learn the task. Throughout training, the rate at which adversarial examples are produced increases by 19.4% while the amount of interaction with the victim model decreases by 53.2%, on average. These results empirically show that agents grow stronger over training by producing more adversarial examples efficiently. We show the generation RL-based adversarial examples in Figure 1.
Next, we analyze how the reward and transition hyperparameters in both RL Max Loss and RL Min Norm affect adversarial examples. The $\epsilon$ parameter controls the amount of distortion the RL Max Loss agent is allowed to have on a given input. The $c$ parameter controls the degree to which the RL Min Norm agent gets rewarded for reducing victim model confidence over minimizing distortion. While training the agent increases the performance of both attacks, we show that it is also dependent on the choice of $\epsilon$ and $c$. For our experiments, we choose $\epsilon=0.5$ and $c=1.0$ according to this sensitivity analysis that balances adversarial objectives. An adversary using these attacks must consider them before attacking to meet their desired goal.
Lastly, we evaluate how the trained agent’s ability to craft adversarial examples generalizes to unseen data and how it performs relative to the widely used optimization-based SquareAttack algorithm (Andriushchenko et al., 2020). On an unseen dataset outside of training, the trained agent’s misclassification rate, median queries, and average distortion across adversarial examples fall in distribution of the training adversarial examples. In a black-box comparison with SquareAttack, we compare $5000$ episodes of the RL attacks against state-of-the-art to show that learning adversarial examples enables the RL attacks to generate 13.1% more adversarial examples. These results highlight the efficacy of an adversary learning adversarial examples through RL over existing methods.
2 Background
Evasion attacks in AML aim to craft adversarial examples: inputs that create a disagreement between human and machine, e.g., making a picture that a human would recognize as bus be classified as an ostrich  (Szegedy et al., 2014). In this the adversary attempts to minimize the perturbation applied to an input to reduce the victim model’s confidence on the true label. Crafting adversarial examples balances (a) minimizing the victim model’s confidence on the true label and (b) minimizing the amount of distortion applied. Here, we describe two classes of algorithms that optimize for one while holding the other to a constraint:
[…]
where $\delta$ is the distortion, $x\in\mathbb{R}^{n}$ the input, $y$ the true label, $\epsilon$ the distortion budget, $L$ the loss function between victim model predictions and true label (e.g., cross-entropy loss), and $Z:\mathbb{R}^{n}\to\mathbb{R}^{m}$ the victim model classifier. Max Loss attacks take steps in the direction of maximizing the victim model loss while enforcing budget constraints by projecting onto the $\ell_{p}$-norm budget ball  (Madry et al., 2019; Andriushchenko et al., 2020). Min Norm attacks optimize a soft constraint optimization objective and use a search algorithm to select the multiplier that enforces the misclassification constraint while taking steps that minimize the $\ell_{p}$-norm of distortion  (Carlini & Wagner, 2017).
Referred to in the security domain as a threat model, solving the optimization depends on access to the victim model $Z$. Here, we define our threat model mode with respect to these classes of attacks:
White-Box Attacks This setting is when the adversary is assumed to have full access to the victim model and its parameters. These attacks are stronger because they can readily apply first and second order gradient methods like projected gradient descent (Max Loss) or stochastic gradient descent on a soft constraint optimization (Min Norm) (Madry et al., 2019; Carlini & Wagner, 2017). These attacks help us understand what the worst-case adversary can do to a machine learning model.
Black-Box Attacks This setting assumes only query-based access to the victim model, i.e., only inputs and outputs. Black-box attacks can be further divided into decision-based and score-based: if the output of the model is a classification (i.e., hard-label) or a vector of class probabilities. Despite limited access, these attacks are effective in achieving their goal (Lucas et al., 2023; Sheatsley et al., 2022; Andriushchenko et al., 2020). While minimizing victim model confidence and distortion are the main AML benchmarks, a third benchmark exists in minimizing victim model queries for black-box attacks. SquareAttack is a score-based black-box attack on images that iteratively crafts random square patches and keeps the patch if it increases the victim model loss (Max Loss) (Andriushchenko et al., 2020). In this paper, we consider a score-based black-box threat model where RL can use victim model confidences for a well-shaped reward.
3 Learning Adversarial Policies
RL enables the adversary to learn from the success and failure of crafting adversarial examples to improve future attacks. Black-box settings involve a feedback loop between the adversary and victim model, allowing us to construct the MDP necessary to use RL. The rest of the section will layout (1) the episodic setting with the MDP, and (2) the attack procedure.
3.1 MDP Formulation
In the adversary’s MDP, each episode begins with a randomly sampled clean input as part of the initial state. During an episode, the adversary applies perturbations to a select set of features and is rewarded in part by the decrease in victim model confidence on the true label of the selected image. Episodes terminate after the victim model is fooled, meaning that an adversarial example has been crafted (or a threshold of steps is reached, and the process has been deemed to fail). Here, we show an example trajectory with a random agent and randomly sampled clean input from CIFAR-10:
Concretely, the processes begin with an input at the start state ($s_{0}$). At each timestep ($s_{1}\dots s_{k}$), the agent applies (human imperceptible) incremental perturbations to the input that lowers the victim model confidence. The episode terminates in state ($s_{T}$) when the model changes its classification to the adversary selected label, or a maximum of steps is reached. Next, we show how we can model this process as an MDP.
3.1.1 States and Actions
The state space of the attack MDP consists of information available to the adversary. In the black-box setting, this includes the victim model outputs on a given sample $Z(\cdot)$ (i.e., oracle). We define the state representation $s_{t}$ of the attack at a given timestep $t$ as:
$s_{t}=(x_{t},y,Z(x_{t}))$ (2)
where $x_{t}\in\mathbb{R}^{n}$ is the resulting input after $t$ steps of the environment starting at input sample $x_{0}$, $y$ the ground-truth label of $x_{0}$, and $Z(x_{t})$ the victim model outputs on $x_{t}$. This information allows the agent to learn effective perturbations at different states of the process.
The actions in the environment represent a distortion to be applied to the current input $x_{t}$. Learning fine-grained distortions on every feature of the input becomes difficult on high-dimensional data (e.g., images, text, or network data). We define the action $a_{t}$ at a given timestep $t$ as a vector of $N$ features and associated distortions $\delta$:
$a_{t}=\left((i_{1},\delta_{1}),(i_{2},\delta_{2}),\dots,(i_{N},\delta_{N})\right)$ (3)
where $i_{1},i_{2},...,i_{N}\in\{1,n\}$ are the $N$ features of the input $x_{t}$ selected. Each selected feature is paired with distortions $\delta_{1},\delta_{2},...,\delta_{N}$ with each having magnitude $|\delta_{j}|\leq\theta$ for $j\in\{1,...,N\}$. This allows us to balance the trade-off between many and few distorted features with large or little distortion through $N$ and $\theta$ for effective and efficient policy learning.
3.1.2 Reward and Transition Functions
As highlighted above, we categorize AML objectives into two classes: Max Loss and Min Norm. Our goal is to engineer reward and transition functions so that RL optimizes these objectives. To quantify the victim model’s confidence of the true label, we use:
$f(x,y)=\max\{[Z(x)]_{y}-\max_{i\neq y}[Z(x)]_{i},0\}$ (4)
where $f$ measure the difference between the victim model’s confidence on the true label $y$ and the highest confidence value that is not of the true label $y$. The function is clipped at 00, which means that any iterate $x_{t}$ that is misclassified by the victim model will have a value $f(x,y)=0$. To model one-step differences in victim model confidence and distortion, let us define:
$\Delta_{t+1}f=f(x_{t},y)-f(x_{t+1},y)\quad;\quad\Delta_{t+1}\delta=\|x_{t}-x_{0}\|_{2}-\|x_{t+1}-x_{0}\|_{2}$ (5)
where $x_{t+1}$ is the input at the next state $s_{t+1}$ and we use the $\ell_{2}$-norm to quantify distortion. For simplicity, we define the function $\phi:\mathbb{R}^{n}\times\mathcal{A}\to\mathbb{R}^{n}$ that takes an input sample $x_{t}\in\mathbb{R}^{n}$ and action $a_{t}\in\mathcal{A}$ and returns the input with specified distortions from the action. The following will describe two versions of an RL attack corresponding to the adversaries’ goals.
RL Max Loss. This setting rewards actions that reduce the victim models confidence. The transition should project the distortion onto a $\epsilon$ distortion budget centered at the starting input $x_{0}$ and keep the distortion if the victim model’s confidence is reduced. Let us define $x_{t}^{a_{t}}=\text{Proj}_{\epsilon}[\phi(x_{t},a_{t})-x_{0}]+x_{0}$ as the candidate input compliant with the $\epsilon$ budget constraint after action $a_{t}$ is applied. Then, the next input $x_{t+1}$ at next state $s_{t+1}$ can be defined with resultant reward $R(s_{t},a_{t})$ as
$x_{t+1}=\big{\{}x_{t}^{a_{t}}\textbf{ if }f(x_{t},y)-f(x_{t}^{a_{t}},y)>0\textbf{ otherwise }x_{t}\big{\}}$ (6)
$R(s_{t},a_{t})=\Delta_{t+1}f$ (7)
where the input at the next state and the reward follow the distortion success in reducing the victim model confidence with the proposed action. It is important to note that a transition involves exactly one query to the victim model to collect outputs $Z(x_{t})$. We denote RL Max Loss transition and reward as $P_{\text{Max Loss}}$ and $R_{\text{Max Loss}}$, respectively.
RL Min Norm. This setting should give rewards to actions that not only reduce victim model confidence but minimize the distortion in doing so. Like RL Max Loss, the state of the next input should result from applying an action that reduces victim model confidence with minimal distortion. Let us define $x_{t}^{a_{t}}=\phi(x_{t},a_{t})$ as the candidate input. The next input $x_{t+1}$ at the next state $s_{t+1}$ can be defined with resultant reward $R(s_{t},a_{t})$ as
$x_{t+1}=\big{\{}x_{t}^{a_{t}}\textbf{ if }c\cdot\big{[}f(x_{t},y)-f(x_{t}^{a_{t}},y)\big{]}+\big{[}\|x_{t}-x_{0}\|_{2}-\|x_{t}^{a_{t}}-x_{0}\|_{2}\big{]}>0\big{\}}$ (8)
$R(s_{t},a_{t})=c\cdot\Delta_{t+1}f+\Delta_{t+1}\delta$ (9)
where $c$ weights the importance of reducing the victim model confidence versus minimizing distortion. We experiment with different values of $c$ in the evaluation to determine a value that balances the effectiveness and efficiency of distortions. Like RL Max Loss, a transition involves one query and the distortion proposed by the action is kept if the change in victim model loss and distortion is positive and rewarded accordingly. We denote RL Min Norm transition and reward as $P_{\text{Min Norm}}$ and $R_{\text{Min Norm}}$, respectively.
3.2 The Reinforcement Learning Attack
We leverage the MDP to launch the RL attack. The attack comes in two forms: (a) attacking while training the agent [^1] Note that in practice, an attacker would train the agent by accessing the model in a subtle way to avoid detection (e.g., by querying the model slowly over time or using known model training data to create a surrogate model to train the agent, (Papernot et al., 2017)). Thereafter, the adversary could use the trained agent to attack the model without restraint. and (b) attacking with a fixed policy. In Figure 2, we detail the attack during the training stage. First, a randomly sampled input and label $(x_{0},y)$ from training dataset $\mathcal{D}$ is used to query the victim model to initialize the start state $s_{0}=(x_{0},y,Z(x_{0}))$. The $(s,a,r,s^{\prime})$ interaction with the victim model proceeds following policy $\pi_{\theta}$ with respect to the attack type RL Max Loss or RL Min Norm. The agent stores interactions and updates the policy according to the specific RL algorithm (on-policy or off-policy). When the episode terminates, the process restarts at start state sampling. After training the agent, the adversary can use the trained adversarial agent to perform the attack in a traditional RL policy evaluation setting. The next section evaluates how well this approach trains adversarial agents that learn better attack strategies.
Figure 2: Training the adversarial agent: (1) a randomly sampled clean input and label $(x_{0},y)$ from training dataset $\mathcal{D}$, (2) the start state $s_{0}=(x_{0},y,Z(x_{0}))$ is initialized with the initial query of $x_{0}$, (3,4) the transition and reward function with respect to RL Max Loss and RL Min Norm mediate interaction with the victim model, (5) the policy $\pi_{\theta}$ is updated according to the RL algorithm.
4 Evaluation
We evaluate the adversarial algorithm by asking the following research questions:
  1. RQ1.
    Is reinforcement learning suitable for this task, and does it improve the effectiveness and efficiency of adversarial example generation?
  2. RQ2.
    How do hyperparameters $\epsilon$ in RL Max Loss and $c$ in RL Min Norm affect adversarial example effectiveness and efficiency?
  3. RQ3.
    Do RL attacks generalize to unseen data and different victim models, and how do they compare to known black-box attacks?
4.1 Experimental Setup
We use the CIFAR-10 image dataset throughout. This dataset is comprised of $70k$ object color images separated into $60k$ training and $10k$ test examples. CIFAR-10 has been widely used to study AML in many models and training architectures. Each RGB image has $3072$ features encoding 32x32 pixel values. For black-box experiments, we use the $10k$ testing examples and create two partitions: $5k$ images for $\mathcal{D}$ in attack training, and $5k$ images as a hidden dataset $\mathcal{D}^{\prime}$ (for policy evaluation). The RL attack training was performed on NVIDIA A100 GPUs with 40 GB of VRAM on pools of the Center for High Throughput Computing (Center for High Throughput Computing, 2006).
Model Architectures. The victim model $Z$ has a ResNet50 (He et al., 2015) architecture implemented in Pytorch. It is trained on the CIFAR-10 training dataset and has a 95.67% accuracy on the testing dataset. Further training details can be found in Appendix A.
For RL training, we use the PPO implementation in StableBaselines3 (Raffin et al., 2021). The adversary’s policy consists of a ResNet18 feature extractor followed by a filly connected feedforward neural network that outputs the parameters of the action distribution. The PPO algorithm also uses a separated value function network which we model as a feedforward network that shares the policy’s ResNet18 feature extractor. PPO hyperparameters and training details are in Appendix A.
Action Hyperparameters. In Section 3.1, we define the actions as selecting $N$ features of the input with associated features $\delta_{i}$ for $i\in\{1,..,N\}$. For our evaluation, we fix $N=3$ and $|\delta_{i}|\leq 0.01$ for all $i\in\{1,2,3\}$ to focus on evaluating the agents ability to learn and generalize adversarial examples.
Metrics. Adversarial examples in all threat models are benchmarked on distortion and victim model misclassification. As such, we use $\ell_{2}$-norm distortion and misclassification (i.e., $\arg\max_{i}Z(x_{T})^{(i)}\neq y$) at terminal states $s_{T}$. The exception appears with black-box threat models, where the number of victim model queries per adversarial example are also considered. Thus, we track the number of victim model queries in each episode. These metrics encapsulate the strength of a black-box adversary and will be used to show their changes as the agent is trained.
4.2 Training Adversarial Agents
In this section we ask RQ1: is reinforcement learning suitable for this task, and does it improve the effectiveness and efficiency of adversarial examples? In other words, is the agent able to learn a policy with greater average return, and does the adversary get stronger with more successful adversarial examples in fewer queries? To answer this, we perform an evaluation in two parts: (1) analyzing the return and adversarial example metrics over training, and (2) policy evaluation on adversarial examples before and after training. We use average return, average $\ell_{2}$-norm distortion, misclassification rate, and median queries of adversarial examples as metrics. We observe that ($a$) if the average return increases, the agent is effectively learning a better policy on the task, ($b$) if the distortion decreases, the agent is learning more effective adversarial examples and ($c$) If median queries decrease, the agent is learning adversarial examples more efficiently.
Training Evaluation In Section 3.1, we define the rewards $R_{\text{Max Loss}}$ and $R_{\text{Min Norm}}$ for the RL Max Loss and RL Min Norm attacks, respectively. Our goal is to show that the average return for each attack increases over training to determine if the task is suitable for RL. Here, we run each attack type for $5000$ episodes against the ResNet50 victim model for 10 trials each and plot the rolling average over the previous $100$ episodes of average return, misclassification rate, median number of queries, and average $\ell_{2}$-norm distortion. We fix $c=1.0$ and $\epsilon=0.5$ for the RL Min Norm and RL Max Loss attacks, respectively. We plot the four metrics with respect to training episode with a 95% confidence interval in Figure 3.
(a) Average Return
(b) Misclassification Rate
(c) Median Queries
(d) Average $\ell_{2}$-norm Distortion
Figure 3: RL Max Loss and RL Min Norm attack training. Average return, misclassification rate, median queries, and average $\ell_{2}$-norm distortion with respect to training episodes for 10 trials per attack with a 95% confidence interval.
We observe that the average return increases over the $5000$ episodes of training in 3(a). This means that the agent is learning a better policy with more discounted reward— an artifact of RL suitability for a given task. The RL Max Loss is offset above the RL Min Norm because the latter penalizes distortion in the reward, whereas the former explicitly projects the distortion onto the specified $\epsilon$ budget.
We also see that the misclassification rate increases by 19.4% in 3(b) and the median number of queries decreases by 53.2% in 3(c), on average. This means that episodes are terminating on victim model misclassifications with higher probability. Further, the agent gets to the misclassifications more efficiently because the median number of queries reduces from beginning to end of training. We see the difference in attack performance in 3(d) where RL Max Loss averages 0.37 $\ell_{2}$-norm distortion throughout training and RL Min Norm increases logarithmically to 0.48 $\ell_{2}$-norm distortion with high variance. This is because the distortion budget is enforced in the former and softly constrained in the latter. Consequently, the misclassification rate in RL Min Norm is 8.3% higher at the end of training. Nevertheless, the results demonstrate RL’s ability to learn adversarial examples.
Policy Evaluation With our black-box evaluation metrics for adversarial examples, we aim to show that adversarial examples produced from a trained policy are more efficient than those produced by a random policy. Here, we take 10 random policies (Random) and 10 trained policies (Trained) for each RL attack after $5000$ episodes of training and evaluate each on 100 images randomly sampled from $\mathcal{D}$. This equates to 1000 episodes of policy evaluation for each RL attack.
In RL Max Loss, the trained agent misclassification rate in the test setting is 75.0% compared to the random agent with 59.1%. Similarly, the RL Min Norm trained agent achieves 86.6% misclassification rate compared to the random agent with 60.7%. We plot the distribution of queries per adversarial example in Figure 4 normalized over the total number of adversarial examples to better understand if the adversarial examples are not only getting more successful but doing so efficiently.
In both attacks, we observe that the adversarial examples crafted with the trained policy are heavily right skewed compared to the random policy. In the RL Max Loss attack distribution (4(a)), we note that the median number of queries for Trained and Random policies are 554 and 1001, respectively. Similarly, in the RL Min Norm attack distribution (4(b)), the median number of queries for Trained and Random policies are 608 and 961, respectively. The steep shift in query distribution of adversarial examples from before training to after training tells us that the RL agents learn an efficient attack strategy to generate more adversarial examples.
(a) RL Max Loss
(b) RL Min Norm
Figure 4: RL Max Loss and RL Min Norm adversarial example query distribution normalized by misclassification rate for trained agents (Trained) and random agents (Random) in each RL attack setting.
4.3 RL Attack Hyperparameter Sensitivity
In Section 3.1, we define two types of RL attacks with shared state and action spaces but different reward and transition functions. Recall that the $\epsilon$ parameter in RL Max Loss controls the maximum distortion allowable, and $c$ in RL Min Norm controls the weight of the reward given to reducing victim model confidence over minimizing distortion. Here, we target RQ2: how do hyperparameters $\epsilon$ in RL Max Loss and $c$ in RL Min Norm affect adversarial example effectiveness and efficiency?
This analysis is broken down into two parts: (1) analyzing the effect $\epsilon$ has on the adversarial examples crafted by RL Max Loss, and (2) analyzing the effect $c$ has on the adversarial examples crafted by RL Min Norm. We use the misclassification rate and $\ell_{2}$-norm distortion on adversarial examples crafted after training to examine how changing $\epsilon$ and $c$ affects adversarial capabilities. We plot the performance of a random agent and trained agent in a policy evaluation setting in Figure 5.
Analyzing the effect of $\epsilon$. With our setup of the RL Max Loss attack, the agent is rewarded by $\Delta_{t+1}f$ at each timestep and all distortions stay within $\epsilon$ $\ell_{2}$-norm distortion budget through $R_{\text{Max Loss}}$ and $P_{\text{Min Norm}}$. We run the RL Max Loss attack for $5000$ episodes with $\epsilon$ from 00 to $1.0$ in $0.1$ increments against the ResNet50 victim model and each tested on 100 images randomly sampled from $\mathcal{D}$. We plot the random and a trained agent misclassification rate with respect to $\epsilon$ in 5(a).
An adversary’s misclassification rate will increase with respect to distortion budget because the adversary is given more space to maneuver. Indeed, we observe that the trained RL Max Loss agent’s performance increases logarithmically with respect to $\epsilon$. We note that because the adversary learns a better policy, the trained agent’s performance tops the random agent by an average of 13.2% across $\epsilon$ values. Depending on the threat model and the adversary’s distortion budget goal, we can observe the performance/budget trade-off when selecting $\epsilon$ to train the agent.
Analyzing the effect of $c$. In our setup of the RL Min Norm attack, the agent is rewarded by $c\cdot\Delta_{t+1}f+\Delta_{t+1}\delta$ at each timestep through $R_{\text{Min Norm}}$ and $P_{\text{Min Norm}}$. We run the RL Min Norm attack for $5000$ episodes with $c$ from 00 to $10^{2}$ in log-scale increments against the ResNet50 victim model and each tested on 100 images randomly sampled from $\mathcal{D}$. We evaluate a random policy before training (Random) and a trained policy (Trained) for 10 trials each and plot the misclassification rate with respect to $\epsilon$ on a 95% confidence interval in 5(b). To reduce legend overload, we denote the trained agent performance as solid lines and random agent performance as dotted.
(a) RL Max Loss ($\epsilon$)
(b) RL Min Norm ($c$)
Figure 5: Hyperparameter sensitivity: (Left) RL Max Loss attack misclassification rate vs. $\epsilon$ distortion budget constraint and (Right) RL Min Norm attack misclassification rate and average $\ell_{2}$-norm distortion vs. $c$. Trained agent (Trained) results are plotted against a random agent (Random) for 10 trials each with a 95% confidence interval.
The adversary’s performance and distortion will increase with respect to $c$ because the agent values decreasing the norm over decreasing victim model confidence. Indeed, we observe this behavior as $c$ increases on log-scale. Further, we see that the agent learns a policy that increases both misclassification rate and distortion. At $c=1.0$, we observe that the trained agent has a misclassification rate 29.8% higher on average and an average $\ell_{2}$-norm distortion 0.8 higher on average. These results are consistent with traditional Min Norm attacks in AML (Carlini & Wagner, 2017). Like our analysis on $\epsilon$ for RL Max Loss, an optimal value of $c$ depends on the adversary’s goals and capabilities.
4.4 RL Attack Generalization and Comparison
Each adversarial agent trains on images from the training dataset $\mathcal{D}$. To investigate a broader landscape of attack capabilities in black-box settings, we target RQ3: do RL attacks generalize to unseen data and different victim models, and how do they compare to known black-box attacks?
We bifurcate the analysis of this question into two experiments: (1) analyzing the adversarial examples crafted on an unseen dataset and (2) comparing the performance of the RL attacks against a known highly-performant and effective black-box attack, SquareAttack (Andriushchenko et al., 2020). The first part will examine adversarial examples crafted after training on randomly selected inputs from the training dataset $\mathcal{D}$ and testing dataset $\mathcal{D}^{\prime}$. The second part will compare the RL attacks against SquareAttack over misclassification rate, median queries, and average distortion.
Generalization Here we evaluate whether the attack generalizes to unseen inputs, i.e., samples not used for training. In Section 4.1, we construct disjoint sets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ for training and testing, respectively, with $5000$ images of support each. After training the attacks, we randomly sample 100 inputs from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ and evaluate the trained policies to record the misclassification, query count, and $\ell_{2}$-norm distortion. We repeat this experiment for 10 trials.
In Table 1, we record the Misclassification Rate (MR), Median Queries (MQ), and Average $\ell_{2}$-norm distortion (Avg. $\ell_{2}$) for each trained RL attack type against train and test datasets $\mathcal{D}$ and $\mathcal{D^{\prime}}$. We see that the adversarial example metrics are very similar when comparing training to testing dataset performance. We perform a two-tailed Mann Whitney U-test to determine if the performance on the test dataset (across each metric) is out-of-distribution of the training dataset. We record the p-values of the hypothesis tests on each metric for each attack and observe that p-values range from 0.20 to 0.92. This means the performance on the test dataset $\mathcal{D}^{\prime}$ is well within the distribution of the train dataset $\mathcal{D}$ performance. Thus, the policy learned by the agent generalizes to unseen inputs.
Dataset RL Max Loss RL Min Norm
MR MQ Avg. $\ell_{2}$ MR MQ Avg. $\ell_{2}$
$(H_{0})$ Train Dataset 0.75$\pm$0.07 549 0.37$\pm$0.02 0.87$\pm$0.05 594 0.41$\pm$0.03
$(H_{1})$ Test Dataset 0.75$\pm$0.06 515 0.36$\pm$0.01 0.85$\pm$0.04 647 0.42$\pm$0.03
p-value ()($\neq$)( ≠ ) 0.92 0.32 0.24 0.20 0.33 0.24
Table 1: Comparison of RL attack types on train and test datasets showing misclassification rate (MR), median queries (MQ), and average $\ell_{2}$-norm distortion (Avg. $\ell_{2}$). The p-value results from the Mann-Whitney U-test that the distribution on the test dataset ($H_{1}$) is not equal ($\neq$) to that of the train dataset ($H_{0}$) for MR, MQ, and Avg. $\ell_{2}$.
Comparison Given the RL attacks’ ability to improve adversarial example performance in black-box settings, we compare the stream of inputs adversarial setting against SquareAttack. Like the MDP setup of the RL agent, we randomly sample an input from dataset $\mathcal{D}$ and apply SquareAttack to record the misclassification, query count, and $\ell_{2}$-norm distortion. For a fair comparison, we denote this as a single episode analogous to the RL attacks. We plot the rolling average misclassification rate, median queries, and $\ell_{2}$ norm distortion for two variations of SquareAttack: $\epsilon\in\{0.4,0.45\}$ against RL Max Loss ($\epsilon=0.5$) and RL Min Norm ($c=1.0$) across each metric in Figure 6.
(a) Misclassification Rate
(b) Median Queries
(c) Average $\ell_{2}$-norm Distortion
Figure 6: SquareAttack ($\epsilon=0.45$), SquareAttack ($\epsilon=0.4$), RL Max Loss, and RL Min Norm attack comparison with respect to episodes for 10 trials per attack with a 95% confidence interval.
On average, the SquareAttack misclassification rate, median queries, and average $\ell_{2}$-norm distortion is (0.691, 784, 0.425) throughout the entire stream of episodes, respectively. For comparison, the RL attacks in the first 100 episodes average (0.656, 1279, 0.361) and in the last 100 episodes average (0.822, 831, 0.411). This shows that by learning adversarial examples through RL, the adversarial agent succeeds 13.1% more after $5000$ consecutive attacks than SquareAttack, shown in 6(a). Further, SquareAttacks constant performance throughout the stream of episodes is an artifact of why an adversary is likely to use an RL approach in a sustained attack setting.
5 Discussion
MDP Improvements The RL attacks are only as good as the action space and reward function that allow the agent to change the state of the input with more reward. The current formulation selects $N$ features of an input with each selected feature getting $\delta_{i},i\in\{1,...,N\}$ with bounded distortion $|\delta_{i}|\leq\theta$ where $\theta\in(0,1)$ for the image domain. Distortion strategies come in many forms in AML, and alternative representations of the action space such as latent-space distortions could enhance the efficiency and attack success in RL. Further, including curriculum learning or adaptive rewards based on attack success could accelerate training. Future work should explore the combinations of different MDP components to develop more practical, real-world adversarial testing.
Cross-Domain applicability The study in this paper focuses on image classifiers, but the framework and analysis is not inherently tied to it. Images provide a natural domain to evaluate new attacks due to well-defined distortion metrics like the $\ell_{2}$-norm and rudimentary constraints the require features to be within $[0,255]$. Nevertheless, these strategies can be readily applied to other classification tasks such as malware and network intrusion detection. Investigating these domains not only studies the domain generalizability of the RL attacks proposed in this paper but also expands their impact on securing a range of machine learning applications.
Relevant connections to RL use in secruity RL has been explored in various security domains, including malware detection (Etter et al., 2023), network intrusion detection (Apruzzese et al., 2020), and image classification (Akers & Barton, 2024; Tsingenopoulos et al., 2019). These works rely on sparse reward functions and discrete action spaces, limiting the adaptability of learned policies. Moreover, these studies primarily evaluate RL-based attacks on a per-sample basis, optimizing perturbations for a fixed input without explicitly considering generalization across multiple examples.
6 Conclusion
This paper explores the capabilities of RL in black-box AML. We developed a novel MDP framework consisting of two agents, RL Max Loss and RL Min Norm, that emulate two classes of AML algorithms. The MDP setup enables the adversary to use RL to learn attack strategies that generate adversarial examples more successfully and with efficiency. Indeed, we find that: (1) the agents learn a policy that improves the misclassification rate by 19.4% and (2) the adversarial examples generated by the agents require 53.2% fewer queries over training on the CIFAR-10 dataset. Further, we analyze the sensitivity of reward hyperparameters $\epsilon$ and $c$ for RL Max Loss and RL Min Norm to show their consistency with state-of-the-art AML algorithms. Last, we demonstrate the trained policy’s ability to generalize crafting adversarial examples to inputs outside of the training dataset and show that in a comparison with a state-of-the-art black-box attack, the RL attacks become 13.1% more successful at generating adversarial examples after $5000$ episodes of training. These findings suggest that adversarial examples can be learned through RL, and that with the improvement of RL as a field results in stronger adversaries.
Broader Impacts
Using RL to craft adversarial examples has significant implications beyond evaluating model robustness. As AI becomes increasingly integrated with healthcare, finance, and cybersecurity, it is imperative that we understand their vulnerabilities and how an adversary can abuse them. RL attacks provide a dynamic way to evaluate models that static methods like traditional AML methods may miss. However, this raises ethical considerations regarding the misuse of these techniques. We emphasize the need for responsible disclosure and robust defense strategies. Future research should balance the offensive advancements with defensive improvements to ensure RL based attacks contributes to better AI security.
Appendix
Appendix A Model Architecture Details
PPO Parameter Value
Feature Extractor ResNet18
Policy Network Linear(128,64,$2N$)
Value Network Linear(128,64,1)
Activation ReLU
Learning Rate (LR) 2.5e-3×$\times\alpha$
$\gamma$ 0.99
GAE $\lambda$ 0.95
Clip Range 0.1×$\times\alpha$
Value Function Coef 0.5
Entropy Coef 0.01
Horizon $T$ 4096
Environments 32
Optimizer Adam
Batch Size 64
Epochs 10
Victim Parameter Value
Architecture ResNet50
Loss Function CrossEntropyLoss
Optimizer SGD
Learning Rate (LR) 0.1
Momentum 0.9
Weight Decay 5e-4
LR Scheduler Cosine Annealing
Batch Size 256
Epochs 50
Test Dataset Accuracy 95.67%
Table 2: PPO algorithm and victim model hyperparameters. $\alpha$ is linearly annealed from 1 to 0 over the course of training in PPO.
Acknowledgments
The authors thank Rahul Chatterjee, Yohan Beugin, and Kunyang Li for their helpful comments in previous iterations of this work.
Funding Acknowledgment. This material is based upon work supported by the National Science Foundation under Grant No. CNS-2343611. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Supplementary Materials
The following content was not necessarily subject to peer review.
Appendix B Action Hyperparameters
Figure 7: Misclassification Rate, Median Queries, and Average $\ell_{2}$-norm Distortion on adversarial examples post-training for different $(N,\theta)$ configurations with RL Min Norm and RL Max Loss attacks.
Appendix C RL Generated Adversarial Examples
Figure 8: CIFAR-10 adversarial examples generated by black-box RL attacks. Each image contains the confidence on the original class and confidence on the incorrect class.